OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (-6, 1) m = 6

OpenStudy (ivettef365):

you have m=6 points (x,y) -> (-6,1) just need to find b, so you arrange it this way y = mx + b 1 = 6(-6) +b solve for b

OpenStudy (anonymous):

i dont know how to solve for b

OpenStudy (ivettef365):

you have 1 = 6(-6) + b 1 = -36 + b +36 +36 37 = b so b= 37 now you have in form of y = mx + b y = 6x + 37